OpenStudy (anonymous):

26 + 6b ³ 2(3b + 4) A. all real numbers B. b £ C. b ³ D. no solutions

OpenStudy (anonymous):

wow i see the problem the inequality you see is being replaced, when you copy and paste, by that little 3

OpenStudy (anonymous):

what do you actually see on the computer?

OpenStudy (anonymous):

I see the little 3 in my question on the test

OpenStudy (anonymous):

hmmm then maybe it is your browser it should be either \(\geq\) or \(\leq \) i think

OpenStudy (anonymous):

maybe you can try a different browser i can tell by the question that they want you to solve an inequality of some kind, but i cannot tell whether it is \(\leq \) or \(\geq\)

OpenStudy (anonymous):

Well skip that one

OpenStudy (anonymous):

you have the little 3 twice, and also that other symbol, so your browser is not reading the script properly it really doesn't make any sense

OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 2|7-7x|=2x+4

OpenStudy (anonymous):

try chrome or firefox

OpenStudy (anonymous):

im on firefox

OpenStudy (anonymous):

\[2|7-7x|=2x+4\] divide both sides by 2 and get \[|7-7x|=x+2\] or to make it easier \[|7x-7|=x+2\]now you have to solve two equations \[7x-7=x+2\] and also \[7x-7=-x-2\] can you solve these?

OpenStudy (anonymous):

No... :/

OpenStudy (anonymous):

really? damn ok lets do this one \[7x-7=x+2\] you want to get \(x\) by itself so subtract \(x\) from both sides and get \[6x-7=2\]then add 7 to both sides and get \[6x=9\] and finally divide both sides by \(6\) to get \[x=\frac{9}{6}=\frac{3}{2}\]

OpenStudy (anonymous):

try this one \[7x-7=-x-2\] this time you want to ADD \(x\) to both sides, then add 7 to both sides, then divide by \(8\)

OpenStudy (anonymous):

x=5/8

OpenStudy (anonymous):

you found it! so the two solutions are \[x=\frac{5}{8}\] or \[x=\frac{3}{2}\]

OpenStudy (anonymous):

:D